OpenStudy (anonymous):

Factor Completely. x^8 - 1/81. ^8 is the exponent of 8. 1/81 is 1 over 81 in a fraction.

OpenStudy (amoodarya):

x^8 - 1/81= x^8 - 1/3^4 =(x^4+1/9)(x^4-1/9) = =(x^4+1/9)(x^2+1/3)(x^2-1/3) =(x^4+1/9)(x^2+1/3)(x+sqrt(1/3))(x-sqrt(1/3))

OpenStudy (amoodarya):